# Assistance-Seeking in Human-Supervised Autonomy:

### v1

@@ Abstract
Using a dual-task paradigm, we explore how robot actions, performance, and the introduction of a secondary task influence human trust and engagement. In our study, a human supervisor simultaneously engages in a target-tracking task while supervising a mobile manipulator performing an object collection task. The robot can either autonomously collect the object or ask for human assistance. The human supervisor also has the choice to rely upon or interrupt the robot. Using data from initial experiments, we model the dynamics of human trust and engagement using a linear dynamical system (LDS). Furthermore, we develop a human action model to define the probability of human reliance on the robot. Our model suggests that participants are more likely to interrupt the robot when their trust and engagement are low during high-complexity collection tasks. Using Model Predictive Control (MPC), we design an optimal assistance-seeking policy. Evaluation experiments demonstrate the superior performance of the MPC policy over the baseline policy for most participants.
@@ I Introduction
With ongoing technological advancements, autonomous systems are increasingly deployed across various sectors including industrial, commercial, healthcare, and agricultural. They offer significant opportunities, such as undertaking hazardous or repetitive tasks and allowing for optimal allocation of human resources. Despite their improving efficiency and reliability, many of these systems still require human supervision and assistance, particularly in dealing with complex tasks and uncertain environments [1, 2, 3, 4]. Autonomous systems are often underutilized by the users due to the lack of trust, defeating the purpose and benefits of using automation. In contrast, some users demonstrate an excessive reliance or trust in automation, which can lead to misuse or abuse of the system[5, 6]. Therefore, for efficient human-robot collaboration, human trust in autonomous systems should be carefully calibrated.
@@
In this work, within a dual-task paradigm involving the supervision of an autonomous assistance-seeking object-collection robot, we study how the robot’s assistance-seeking affects human trust in the robot and their engagement in the secondary task. We also look into how robot performance impacts supervisors’ secondary task performance. We model the dynamics of trust in the robot and engagement in the secondary task using LDS, which are then utilized to design an optimal assistance-seeking policy.
@@
To gain insight into human trust, it is important to understand the various factors that shape it. The effect of robot reliability, workload, and information quantity on human trust has been studied [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Additionally, environmental factors such as task complexity affect human trust as well. A comprehensive investigation into factors affecting human trust in automation categorizes them into three main groups: robot-related, human-related, and environment-related factors [17].
@@
Alongside the study of factors influencing trust, several models of trust evolution have been proposed, and they are categorized into various groups. A class of these models is called probabilistic models. Probabilistic trust models treat trust as a latent state, estimating its distribution within a Bayesian framework [18]. This distribution is conditioned on factors such as human actions, robot performance, and other contextual information relevant to the task. A subclass of such models is POMDP-based models, where the dynamics of trust are defined by the state transition functions conditioned on actions and contextual information, with observations treated as human actions [14, 15, 19, 16, 20, 21, 22]. Such models have been applied for planning to minimize human interruption [14, 22]. These models have also been used in tasks where it is needed to optimize the delivery of information to humans, such that efficiency in the collaboration is maintained [16, 20, 21].
@@
Deterministic linear dynamical systems are used to represent trust dynamics, with trust level, cumulative trust level, and expectation bias as system states. The inputs to these models include factors that influence human trust, such as robot performance and task-specific contextual information [11, 12, 13]. Additionally, alongside human trust towards automation, models have been developed to capture automation’s trust towards humans [23, 24]. Process and measurement noises have also been incorporated into linear models of human trust. To account for these noises, filtering techniques are necessary to accurately estimate the state using output feedback information [25]. In this work, we focus on a similar noisy linear model of trust. In contrast to [25] which uses the traditional linear Kalman filter for trust estimation, our nonlinear measurement model necessitates the use of nonlinear filtering techniques, specifically the particle filter. Furthermore, we leveraged our learned model to determine and experimentally evaluate an optimal assistance-seeking policy.
@@
Using a dual-task paradigm, we investigate how robot actions and performance impact human trust and secondary task engagement. Human participants perform both supervisory and target-tracking tasks simultaneously.We explore an assistance-seeking policy for a robot as a function of human trust, secondary task engagement, and task-specific contexts. We conduct two sets of experiments with human participants. The initial data collection experiment involves a randomized assistance-seeking policy for the robot and random task complexity assignments. Using the collected data, we estimate models for trust dynamics and secondary task engagement using a noisy linear dynamical system. In this model, the dynamics of human trust and secondary task engagement are influenced by the actions of the robot and human, the outcomes of those actions, and the complexities of both tasks. Additionally, we develop a human action model to define the action selection strategy of the human supervisor. Using the estimated models, we apply Model Predictive Control (MPC)[26] to identify the optimal actions based on the current states and task complexities. Subsequently, in the second set of experiments, we evaluate the MPC-based assistance-seeking policy and compare it with a baseline policy. The contributions of this work are:
@@
The rest of the paper is organized as follows. The dual-task setup and tasks are described in Section II. We describe the human behavior model in Section III. We describe the design of MPC policy and its evaluation using a second set of experiments in Section IV. We then summarise and conclude in Section V.
@@ II Dual Task Setup: Robot Supervision and target-tracking task
In our experiment, we rely on the dual-task paradigm, wherein the human participant engages in two simultaneous tasks. Specifically, the human participant supervises an autonomous mobile manipulator tasked with retrieving objects from shelves, while simultaneously performing a target-tracking task. The experiment interface is shown in Fig. 1. We now define both tasks in more detail.
@@ II-B Target-tracking Task
The human participants are instructed to track the ball (target) for a total cumulative time of 5 seconds. Their target-tracking performance $p$ is determined by taking the goal time of 5 seconds and dividing it by the total time it took to finish the trial. The participants were instructed to achieve a performance of at least $0.75$. Then, the reward for the target-tracking task is defined by
@@ II-C Data Collection for Modeling Human Behavior
To build a model of human supervisory behavior, we conducted a study 111 The human behavioral experiments were approved under Michigan State University Institutional Review Board Study ID 391. with $11$ human participants performing the above supervisory and target-tracking tasks simultaneously.
For the supervisory object collection tasks, each participant performed 30 low-complexity and 30 high-complexity trials. An initial set of 60 such trials was selected and was randomly permuted independently for each participant. When the manipulator operates autonomously, it has a success probability of $p^{\textup{suc}}_{H}=0.75$ in high complexity environments and $p^{\textup{suc}}_{L}=0.96$ in low complexity environments. In each trial, the robot asks for human assistance with a probability of 0.3 in high-complexity trials, and with a probability of 0.1 in low-complexity trials. After each trial, participants reported their trust in the autonomous mobile manipulator through an interface with an 11-point scale. For each object collection trial, the data collected are the complexity $C^{1}$, robot action $a^{R}$, human action $a^{H}$, outcome $E$, and the reported human trust $y^{T}$.
For the target-tracking task, participants performed 30 slow-speed trials and 30 normal-speed trials. An initial set of 60 such trials was selected and was randomly permuted independently for each participant and independently of the object-collection task. In each trial, the data collected are the complexity of the trial $C^{2}$, and the performance $p$.
@@ III Human Behavior Model
In the following, we denote trust at the beginning of trial $t$ by $T_{t}\in\mathbb{R}$, the target-tracking engagement in trial $(t-1)$ by $P_{t}\in[0,1]$, the complexities of object collection and target-tracking task by $C^{1}_{t}$ and $C^{2}_{t}$, respectively, the human and robot actions during collaborative object collection in trial $t$ by $a^{H}_{t}$ and $a^{R}_{t}$, respectively, and the outcome of the object collection in trial $t$ by $E_{t+1}$.
@@ III-A Human Trust Dynamics
Let $u^{T}_{t}\in\{0,1\}^{7}$ be an indicator vector that represents one of the above 7 possible events in collaborative object collection trial $t$. We model the trust dynamics using the following Linear Dynamical System (LDS)
@@
$A^{T}$ $=0.92,$
$B^{T}$ $=\begin{bmatrix}0.76&-0.38&0.26&0.78&-0.43&0.52&-0.12\end{bmatrix},$
$\sigma^{2}_{T}$ $=0.22,$
$C^{T}$ $=1.00,$
$\sigma^{2}_{y}$ $=0.22.$
@@ III-B Human Target-tracking Engagement Dynamics
$\phi^{2}_{t}$ $(C^{\textup{2,norm}}_{t},a^{R-}_{t},E^{+}_{t}$ $\phi^{6}_{t}$ $(C^{\textup{2,norm}}_{t},a^{R+}_{t},E^{+}_{t})$
@@
Let $u^{P}_{t}\in\{0,1\}^{8}$ be an indicator vector that represents one of the above 8 possible events in target-tracking task during trial $t$. We model the human target-tracking engagement dynamics as an LDS
$P_{t+1}$ $=A^{P}P_{t}+B^{P}u^{P}_{t}+v^{P}_{t},$ (3)
$p_{t}$ $=C^{P}P_{t}+w^{P}_{t},$ (4)
where $A^{P}\in[0,1]$, $B^{P}\in[0,10]^{1\times 8}$, $v^{P}_{t}\sim\mathcal{N}(0,\,\sigma^{2}_{P})\,,t\in\mathbb{N}$ are i.i.d. realizations of zero-mean Gaussian noise, $C^{P}\in[0,1]$, and $w^{P}_{t}\sim\mathcal{N}(0,\,\sigma^{2}_{p})\,,t\in\mathbb{N}$ are i.i.d. realizations of zero-mean Gaussian noise. The entries of $B^{P}$ determine the influence of the events $\phi^{i}_{t}$ on the human engagement in the tracking task.
The engagement model parameters include $A^{P},B^{P},C^{P},\sigma^{2}_{P}$ and $\sigma^{2}_{p}$. Considering $p$ as the tracking task result of the human, we adopt the EM algorithm to estimate these parameters. The estimated parameters of the engagement dynamics model are
$A^{P}$ $=0.19,$
$B^{P}$ $=\begin{bmatrix}7.47&6.72&7.24&6.38&7.30&6.51&7.06&6.59\end{bmatrix},$
$\sigma^{2}_{P}$ $=1.44,$
$C^{P}$ $=9.96,$
$\sigma^{2}_{p}$ $=3.79.$
The elements of $B^{P}$ represent the effect of each scenario on human tracking task engagement and performance.
The estimated $B^{P}$ indicates that whenever the robot asks for assistance, for the majority of cases, the associated estimated values are greater than the values associated with the cases when the robot collects autonomously. This is expected as whenever the robot asks for assistance, the human supervisor can focus more on their tracking task and finish it before teleoperating the robot, knowing that the robot will wait for them to finish the tracking task. In contrast, when the robot attempts to collect autonomously, the human supervisor deviates their attention from the tracking task to the supervisory task to decide whether to rely or not. The estimated parameters are higher when the target speed is slow compared to when it is normal. Whenever the previous experience with the robot is $E^{-}$, the parameters are lower compared to the case of $E^{+}$. This is consistent with the fact that whenever humans perceive a failure, they tend to pay more attention to the supervisory task in the future.
@@ III-C Human Action Model
We now focus on modeling the probability of the human relying on the robot when the robot attempts to collect autonomously. We assume that this probability depends on the object collection task complexity $C^{1}_{t}$, human trust $T_{t}$, and human target-tracking engagement $P_{t}$, which is representative of the attention they put on the supervisory task. We model the human action probability as a sigmoid function
$\mathbb{P}(a^{H+}|T,P,C^{1})=\frac{1}{1+e^{-(a_{T}^{C^{1}}T+a_{P}^{C^{1}}P+b^{C^{1}})}}.$ (5)
@@
The probabilities of reliance for each object collection task complexity are shown in Fig. 2. For low complexity $C^{1,L}$, the probability of reliance is high and close to 1, regardless of $T$ and $P$. For high complexity $C^{1,H}$, there is a lower probability of reliance in general as compared to low complexity. It can be seen that the probability of reliance increases with trust as well as with tracking task engagement. The latter may be caused by the human focusing more on the tracking task as compared to the supervisory task, thereby increasing their probability of reliance on the mobile manipulator.
Figure 2: Human action model: Probability of reliance $\mathbb{P}(a^{H+}|T,P,C^{1},a^{R+})$ conditioned on object collection task complexity $C^{1}$,$T$, and $P$
Fig.  3 illustrates the relationship between the exogenous factors, trust dynamics, tracking task engagement dynamics, and human action model.

### v2

@@ Abstract
Using a dual-task paradigm, we explore how robot actions, performance, and the introduction of a secondary task influence human trust and engagement. In our study, a human supervisor simultaneously engages in a target-tracking task while supervising a mobile manipulator performing an object collection task. The robot can either autonomously collect the object or ask for human assistance. The human supervisor also has the choice to rely on or interrupt the robot. Using data from initial experiments, we model the dynamics of human trust and engagement using a linear dynamical system (LDS). Furthermore, we develop a human action model to define the probability of human reliance on the robot. Our model suggests that participants are more likely to interrupt the robot when their trust and engagement are low during high-complexity collection tasks. Using Model Predictive Control (MPC), we design an optimal assistance-seeking policy. Evaluation experiments demonstrate the superior performance of the MPC policy over the baseline policy for most participants.
@@ I Introduction
Technological advancements have led to the widespread deployment of autonomous systems in industrial, commercial, healthcare, and agricultural sectors. These systems can handle hazardous or repetitive tasks, optimizing human resource allocation. However, despite their increasing efficiency and reliability, they often still need human supervision for complex tasks and in uncertain environments  [1, 2, 3, 4].
Trust in human-robot interaction is formally defined as “the attitude that an agent will help achieve an individual’s goals in a situation characterized by uncertainty and vulnerability” [5]. Autonomous systems are often underutilized due to the lack of trust, defeating the purpose and benefits of using automation. In contrast, an excessive reliance or trust in automation can lead to misuse or abuse of the system[6, 7]. Therefore, for efficient human-robot collaboration, human trust in autonomous systems should be carefully calibrated.
Supervisors often juggle multiple tasks, managing their own responsibilities while overseeing others. This dynamic also applies to human supervisors overseeing autonomous agents. Supervisors must maintain their task performance even while managing these agents. For instance, imagine a human and an autonomous robot working together in an orchard, collecting fruits. The human is tasked with both collecting fruits and ensuring the robot functions correctly. The human should only intervene with the robot when absolutely necessary to prevent compromising their own fruit collection. Similarly, the robot should be designed to operate with minimal interference to the human’s tasks.
@@
In this work, within a dual-task paradigm involving the supervision of an autonomous assistance-seeking object-collection robot, we study how the robot’s assistance-seeking affects human trust in the robot and their engagement in the secondary task. We also investigate how robot’s actions and performance impact supervisors’ secondary task performance. We model the dynamics of human trust in the robot and secondary-task engagement using LDS, which are then utilized to design an optimal assistance-seeking policy.
@@
To gain insight into human trust, it is important to understand the various factors that shape it. The effect of robot reliability, workload, self-confidence, and information transparency on human trust has been studied [8, 9, 10, 11, 12, 13, 14, 15, 16, 2]. These works highlight that negative experiences lead to a decrease in human trust. Human reliance on automation is influenced not only by trust but also by self-confidence. When individuals have lower self-confidence than their trust in automation, they are more likely to rely on automated systems [9]. Additionally, a higher level of transparency, i.e., providing more information to the human, can often have a positive effect on human trust but can increase their workload [13, 2]. Environmental factors such as task complexity affect human trust as well. A comprehensive investigation into factors affecting human trust in automation categorizes them into three main groups: robot-related, human-related, and environment-related factors [17], and suggests that robot-related factors such as reliability have the highest effect on trust, followed by environment-related factors.
@@
Alongside the study of factors influencing trust, several models of trust evolution have been proposed, and they are categorized into various groups. A class of these models is called probabilistic models. Probabilistic trust models treat trust as a latent state and estimate its distribution within a Bayesian framework [18]. This distribution is conditioned on factors such as human actions, robot performance, and other contextual information relevant to the task. A subclass of such models is Partially Observable Markov Decision Process (POMDP)-based models, where the dynamics of trust are defined by the state transition functions conditioned on actions and contextual information, with observations treated as human actions [15, 16, 19, 2, 20, 21, 22]. Such models have been applied for planning to minimize human interruption [15, 22]. These models have also been used in tasks where it is needed to optimize the amount of information provided to humans, such that efficiency in the collaboration is maintained [2, 20, 21].
@@
Deterministic linear dynamical systems are used to represent trust dynamics, with trust level, cumulative trust level, and expectation bias as system states. The inputs to these models include factors that influence human trust, such as robot performance and task-specific contextual information [12, 13, 14]. Additionally, alongside human trust towards automation, models have been developed to capture automation’s trust towards humans [23, 24]. Process and measurement noises have also been incorporated into linear models of human trust. To account for these noises, filtering techniques are necessary for state estimation using output measurements [25]. Similarly to [25], we focus on a noisy linear trust model; however, our measurement model is nonlinear (binary). The latter necessitates nonlinear filtering techniques, specifically the particle filter.
Additionally, we leverage our learned model to determine and experimentally evaluate an optimal assistance-seeking policy.
@@
Using a dual-task paradigm, we investigate how robot actions and performance impact human trust and secondary task engagement. Human participants perform both supervisory and target-tracking tasks simultaneously. We explore an assistance-seeking policy for a robot as a function of human trust, secondary task engagement, and task-specific contexts. We conduct two sets of experiments with human participants. The initial data collection experiment involves a randomized assistance-seeking policy for the robot and random task complexity assignments. Using the collected data, we estimate models for trust dynamics and secondary task engagement using an LDS. In this model, the dynamics of human trust and secondary task engagement are influenced by the robot and human actions, the outcomes of those actions, and the complexities of both tasks. Additionally, we develop a model to predict human action as a function of their trust and engagement. We leverage Model Predictive Control (MPC)[26] with the estimated models to compute the robot actions to maximize team performance. Subsequently, we evaluate the MPC-based assistance-seeking policy in the second set of experiments and compare it with a baseline policy. The contributions of this work are:
@@
The rest of the paper is organized as follows. The dual-task setup and tasks are described in Section II. We describe the human behavior model in Section III. We describe the design of MPC policy and its evaluation using a second set of experiments in Section IV. We discuss our findings and connect them to a broader context in Section V and and conclude in Section VI.
@@ II Dual Task Setup: Robot Supervision and target-tracking task
In our experiment, we rely on the dual-task paradigm, wherein the human participant engages in two simultaneous tasks. The human participant supervises an autonomous mobile manipulator tasked with retrieving objects from shelves, while simultaneously performing a target-tracking task. The experiment interface is shown in Fig. 1. We now define both tasks in more detail.
@@ II-B Target-tracking Task
The human participants are instructed to track the ball (target) for a total cumulative time of 5 seconds. Their target-tracking performance $p$ is determined by taking the goal time of 5 seconds and dividing it by the total time it took to finish the trial. The participants were instructed to achieve a performance of at least $0.75$.
@@ II-C Data Collection for Modeling Human Behavior
To build a model of human supervisory behavior, we conducted a study [^1] The human behavioral experiments were approved under Michigan State University Institutional Review Board Study ID 9452. with $11$ human participants ($4$ males and $7$ females), performing the above supervisory and target-tracking tasks simultaneously. Participants were recruited via e-mail for the in-person experiment. Participants were first and second-year college students not in the engineering program. An experiment’s total duration was one hour and participants were compensated \$$\$15$ for their participation.
For the supervisory object collection tasks, each participant performed $30$ low-complexity and $30$ high-complexity trials, randomly distributed throughout the experiment. When the manipulator operates autonomously, it has a success probability of $p^{\textup{suc}}_{H}=0.75$ in $C^{\textup{1,H}}$ and $p^{\textup{suc}}_{L}=0.96$ in $C^{\textup{1,L}}$. In each trial, the robot asks for human assistance with a probability of 0.3 in $C^{\textup{1,H}}$ trials, and with a probability of 0.1 in $C^{\textup{1,L}}$ trials. After each trial, participants reported their trust in the autonomous mobile manipulator through an interface with an 11-point scale. For each object collection trial, the data collected are the complexity $C^{1}$, robot action $a^{R}$, human action $a^{H}$, outcome $E$, and the reported human trust $y^{T}$.
For the target-tracking task, participants performed $30$ slow-speed trials and $30$ normal-speed trials. To ensure that all participants do the same number of trials per complexity, an initial set of $60$ such trials was selected and trials were randomly permuted for each participant and independently of the object-collection task. In each trial, the data collected are the complexity of the trial $C^{2}$, and the performance $p$.
II-D Data Summary
For the supervisory task, after a successful autonomous collection, the mean trust ratings of participants are $7.38$ for low complexity and $7.36$ for high complexity, with an overall mean trust rating of $7.37$ across both complexities. After a failure, the mean trust ratings were $6.36$ for low complexity and $6.68$ for high complexity, with an overall mean of $6.6$ across both complexities. When the robot asked for assistance, the mean trust ratings were $7.3$ for low complexity and $6.98$ for high complexity, averaging $7.03$ for both complexities combined. However, when the human voluntarily interrupted, the mean trust rating dropped to $6.29$ when they interrupted the robot in high-complexity, $5$ in low-complexity, and $6.17$ across both complexities.
For the tracking task, at slow speeds, the mean performance was $0.89$ with $90\%$ of the trials having $p\geq 0.75$. At normal speeds, the mean performance dropped to $0.82$ with $78\%$ of the trials having $p\geq 0.75$. As expected, $p<0.75$ whenever participants interrupted the robot.
@@ III Human Behavior Model
In the following, we denote trust at the beginning of trial $t$ by $T_{t}\in\mathbb{R}$, the target-tracking engagement in trial $(t-1)$ by $G_{t}\in[0,1]$, the complexities of object collection and target-tracking task by $C^{1}_{t}$ and $C^{2}_{t}$, respectively, the human and robot actions during collaborative object collection in trial $t$ by $a^{H}_{t}$ and $a^{R}_{t}$, respectively, and the outcome of the object collection in trial $t$ by $E_{t+1}$.
@@ III-A Human Trust Dynamics
Let $u^{T}_{t}\in\{0,1\}^{7}$ be an indicator vector representing one of the above 7 possible events in collaborative object collection trial $t$. We model the trust dynamics using the following LDS
@@
$A^{T}$ $=0.92,\quad\sigma^{2}_{T}=0.22,\quad C^{T}=1.00,\quad\sigma^{2}_{y}=0.22,$
$B^{T}$ $=\begin{bmatrix}0.76&-0.38&0.26&0.78&-0.43&0.52&-0.12\end{bmatrix}.$
@@ III-B Human Target-tracking Engagement Dynamics
$\phi^{2}_{t}$ $(C^{\textup{2,norm}}_{t},a^{R-}_{t},E^{+}_{t})$ $\phi^{6}_{t}$ $(C^{\textup{2,norm}}_{t},a^{R+}_{t},E^{+}_{t})$
@@
Let $u^{G}_{t}\in\{0,1\}^{8}$ be an indicator vector representing one of the above 8 possible events in the target-tracking task during trial $t$. We model the human target-tracking engagement dynamics as an LDS
$G_{t+1}$ $=A^{G}G_{t}+B^{G}u^{G}_{t}+v^{G}_{t},$ (3)
$p_{t}$ $=C^{G}G_{t}+w^{G}_{t},$ (4)
where $A^{G}\in[0,1]$, $B^{G}\in[0,10]^{1\times 8}$, $v^{G}_{t}\sim\mathcal{N}(0,\,\sigma^{2}_{G})\,,t\in\mathbb{N}$ are i.i.d. realizations of zero-mean Gaussian noise, $C^{G}\in[0,1]$, and $w^{G}_{t}\sim\mathcal{N}(0,\,\sigma^{2}_{p})\,,t\in\mathbb{N}$ are i.i.d. realizations of zero-mean Gaussian noise. The entries of $B^{G}$ determine the influence of the events $\phi^{i}_{t}$ on the human engagement in the tracking task.
The engagement model parameters include $A^{G},B^{G},C^{G},\sigma^{2}_{G}$ and $\sigma^{2}_{p}$. Considering $p$ as the tracking task result of the human, we adopt the EM algorithm to estimate these parameters.
$A^{G}$ $=0.19,\quad\sigma^{2}_{G}=1.44,\quad C^{G}=9.96,\quad\sigma^{2}_{p}=3.79,$
$B^{G}$ $=\begin{bmatrix}7.47&6.72&7.24&6.38&7.30&6.51&7.06&6.59\end{bmatrix}.$
The elements of $B^{G}$ represent the effect of each scenario on human tracking-task engagement and performance.
The estimated $B^{G}$ indicates that whenever the robot asks for assistance, for the majority of cases, the associated estimated values are greater than the values associated with the cases when the robot collects autonomously. This is expected as whenever the robot asks for assistance, the human supervisor can focus more on their tracking task and finish it before teleoperating the robot, knowing that the robot has stopped and will not move unless they take control of it. In contrast, when the robot attempts to collect autonomously, the human supervisor may switch attention from the tracking task to the supervisory task to decide whether to rely on the robot. The estimated parameters are higher when the target speed is slow compared to when it is normal. Whenever the previous experience with the robot is $E^{-}$, the parameters are lower compared to the case of $E^{+}$. This is consistent with the fact that whenever humans perceive a failure, they tend to pay more attention to the supervisory task in the future.
@@ III-C Human Action Model
We now focus on modeling the probability of the human relying on the robot when the robot attempts to collect autonomously. We assume that this probability depends on the object collection task complexity $C^{1}_{t}$, human trust $T_{t}$, and human target-tracking engagement $G_{t}$, which is representative of the attention they put on the supervisory task. We model the human action probability as a sigmoid function
$\mathbb{P}(a^{H+}|T,G,C^{1})=\frac{1}{1+e^{-(a_{T}^{C^{1}}T+a_{G}^{C^{1}}G+b^{C^{1}})}}.$ (5)
@@
The probabilities of reliance for each object collection task complexity are shown in Fig. 2. For low complexity $C^{1,L}$, the probability of reliance is high and close to 1, regardless of $T$ and $G$. For high complexity $C^{1,H}$, there is a lower probability of reliance in general as compared to low complexity $C^{1,L}$. It can be seen that the probability of reliance increases with trust as well as with tracking task engagement. The latter may be caused by the human focusing more on the tracking task as compared to the supervisory task, thereby increasing their probability of reliance on the mobile manipulator.
Figure 2: Human action model: Probability of reliance $\mathbb{P}(a^{H+}|T,G,C^{1},a^{R+})$ conditioned on object collection task complexity $C^{1}$,$T$, and $F$
Fig. 3 illustrates the relationship between the exogenous factors, trust dynamics, tracking task engagement dynamics, and human action model.